The sum of one-fifth of Lucy's age and one-seventh of her age is 12. Use algebra to workout her age. Plz help
Answer by jorel555(1290)

Let n be Lucy's age. Then
n/5+n/7=12
7n+5n=420
n=420/12=35 ☺☺☺